# Specific heat of Gd3+- and Eu2+-based magnetic compounds

### v1

@@ IV Effect of anisotropy
In this Section we address the role of the anisotropy first in mean field, and later we incorporate the effect of quantum fluctuations.
@@ IV.1 Mean-field approximation
In this subsection, we discuss the effect of the term $H_{K}=K[3J^{2}_{z}-J(J+1)]$ on the specific heat calculated in the mean-field approximation, in which it is assumed that each site is subjected to a constant effective magnetic field that results from the interaction with the remaining sites. We assume that all sites are equivalent and there is no external magnetic field. For simplicity we restrict the analysis to the easy axis case $K<0$ for which the spins point in the direction of the anisotropy. In the next subsection, both signs of $K$ and the effects of quantum fluctuations will be considered.
@@
$\tilde{B}=g\mu_{B}B_{\text{eff}}=I\langle J_{z}\rangle\text{,~{}~{}~{} }\langle J_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{Me^{-\beta E_{M}}}{Z}\text{,~{}~{}~{} }Z=\sum\limits_{M=-J}^{J}e^{-\beta E_{M}},$ (3)
where $g$ is the gyromagnetic factor, $\mu_{B}$ is the Bohr magneton, $I=\Sigma_{\delta}z_{\delta}J_{\delta}$ is the sum of all interactions ($z_{\delta}$ is the number of neighbors at distance $\delta$ and $J_{\delta}$ its intensity), $\beta=1/(k_{B}T)$ where $T$ is the temperature and $k_{B}$ is the Boltzmann constant. The energies $E_{M}$ for each eigenvalue $M$ of the operator $J_{z}$ is given by
$E_{M}=-\tilde{B}M+K[3M^{2}-J(J+1)].$ (4)
Eqs. (3) and (4) permit to determine $\tilde{B}$ selfconsistently. In the limit $\tilde{B}\rightarrow 0$ one obtains an equation for the critical temperature
$k_{B}T_{C}^{\text{ MF}}=I\langle J^{2}_{z}\rangle\text{, with~{}~{}}\langle J^{2}_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{M^{2}e^{-\beta E_{M}}}{Z}.$ (5)
The specific heat is obtained from the numerical derivative of the entropy $S$ with respect to the temperature
@@
In Fig. 3 we show the resulting effect of the axial anisotropy on the specific heat keeping the value of $k_{B}T_{C}=1$ constant, as the unit of energy. For $K=0$, the value of the intensity for a given $T_{C}$ is $I=3k_{B}T_{C}/[J(J+1)]$, leading to $I=4/21\sim 0.190$ for $k_{B}T_{C}=1$.
For $K=-0.02$, we obtain $I=0.1531$ from Eq. (5). This is because the expectation value $\langle J^{2}_{z}\rangle$ that enters Eq. (5) is larger for $K<0$ and should be compensated with a smaller $I$. In addition, since the values of high $J_{z}$ projection $|M|$ are favored, the effective magnetic field is more efficient than for $K=0$ in splitting the lowest energy levels and therefore the entropy and the specific heat decrease for $T<0.4T_{C}$ with respect to the isotropic case. This is compensated by an increase in $C/T$ for $0.5T_{C}<T<T_{C}$.
@@
IV.2 Beyond mean field
In this Section we discuss the effect of anisotropy in cases where the fluctuations of the effective field are also included.
Figure 4: (Color online) Specific heat over temperature as a function of temperature for several anisotropies for the simple cubic structure including 8 NN and 6 fourth NN with $J_{\delta}=-1$.
In Fig. 4 we show how the specific heat for a structure with $z=14$ is modified by the addition of $H_{K}$. The results are qualitatively the same as in the mean-field case for $T<T_{C}$. Addition of the anisotropy term shifts the entropy from the region $T<0.5T_{C}$ mainly to the region $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
However, in contrast to the mean-field case, addition of the anisotropy term decreases the specific heat for $T>T_{C}$. This reduction is expected because the specific heat in this region primarily results from quantum fluctuations, which are diminished by the energy level splitting induced by $H_{K}$.
@@ V Comparison of different structures with the same effective number of neighbors
In Section III we have discussed the effect of increasing number of neighbors $z$ in the specific heat keeping for simplicity $J_{\delta}=J_{1}$. Here we investigate the effect of changing $J_{\delta}$ for first, second and possibly third NN, keeping the same $r$ [see Eq. (2)], which corresponds to the inverse of the effective number of neighbors $z$. In Fig. 5 (b) we show the comparison for a moderately small $z$: a cubic bcc structure with 8 NN and 6 fourth NN. We compare that case of NN interaction $J_{1}=-1$ and NNN interaction $J_{2}=0$ with $J_{1}=-1/8$ and $J_{2}=-1$. In spite of the radical change in the magnitude of the interactions, the specific heats are rather similar.
In Fig. 5 (c) we show a similar comparison for an fcc structure with different combinations of $J_{1}$ and $J_{2}$. Again the results are similar, except for the case in which the dominant interaction is changed drastically from NN to NNN. When the dominant interaction is that of NNN, the specific heat is larger and the difference is of the order of that between different $z$ for $J_{1}=-1$ (Fig. 5 (a)). Finally in (Fig. 5 (d)) we show a similar comparison for a larger $z$. In this case, the agreement between the two curves is excellent.
@@ VI Fits of experimental data
In this Section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. Previously, the specific heat of GdCoIn5 has been fitted Facio et al. (2015).
The fitting parameters are the known $T_{C}$, which determines the magnitude of the interactions $J_{\delta}$ assumed equal, the number of neighbors $z$ and for the Eu compounds, the anisotropy $K$. The Gd compounds could be fitted with $K=0$. The results are presented in Fig. 6. The Gd and Eu compounds are ordered with increasing effective number of neighbors $z$ in the fit.
@@
GdNi3Ga9 Silva (2017); Nakamura et al. (2023) clearly correspond to a low coordination number ($z=4$). The honeycomb lattice described in Section III with antiferromagnetic interactions is similar to the real structure of the material and provides a very good fit. For the rest of the structures, ferromagnetic interactions are assumed.
GdPdBi Jesus et al. (2014) is a half-Heusler structure with antiferromagnetic coupling. We obtain a very good fit assuming also a layered honeycomb lattice as for the previous compound, but with larger number of effective neighbors (10).
@@
GdCu2Ge2 Duong (2002) has a BCC structure. Using only the 8 NN interactions, the resulting specific heat is very near the experimental one, although slightly below (above) the experimental one for $T<T_{C}$ ($T>T_{C}$) suggesting that $z$ might be a little bit larger.
@@
For GdNiSi3 Arantes et al. (2018) a simple cubic structure with $z=26$ including neighbors up to the 3rd coordination shell provides a very good fit of the data. This completes the Gd compounds that we have considered.
@@
In previous work Facio et al. (2015), the specific heat of GdCoIn5 has been fitted. The compound has a tetragonal structure (near to the simple cubic one). The authors have extrapolated to the thermodynamic limit, the specific heat that corresponds to $z=6$ NN antiferromagnetic interactions, providing an excellent fit of the data.
@@
Now turning to Eu compounds, the specific heat of Eu2Pd3Sn I. Čurlík , is very well fitted using a simple cubic structure with 6 NN and 8 fourth NN ($z=14$) with $J_{1}=J_{2}$ and $K=-0.4J_{1}$. Small deviations exist at very low temperatures, where our numerical method is not fully reliable.
@@
Finally, the same simple cubic structure, but with $K=-0.6J_{1}$ provides a very good fit of the specific heat of EuPdSn2 Čurlík et al. (2018). For temperatures very near $T_{C}$, the theoretical data are a little above (below) the experimental ones for $T<T_{C}$ ($T>T_{C}$)
@@
Figure 6: (Color online) Specific heat over temperature as a function of temperature for several compounds and our fit. $K=0$ for the Gd compounds. Experimental data were taken from Refs. Silva, 2017; Nakamura et al., 2023; Jesus et al., 2014; Duong, 2002; Arantes et al., 2018; Čurlík et al., 2018; I. Čurlík, .
@@ VII Summary
Using quantum Monte Carlo, we have calculated the specific heat of a large number of non-frustrated magnetic structures. As expected from mean-field theory, the critical temperature $T_{C}$ increases approximately proportionally to $I=\Sigma_{\delta}z_{\delta}J_{\delta}$, where $J_{\delta}$ is exchange interaction at distance $\delta$, and $z_{\delta}$ is the number of neighbors at this distance. However $T_{C}$ lies below the mean-field value.
Rather surprisingly, we find that the specific heat depends mainly on the effective number of neighbors $z$ [see Eq. (2)] and is not quite sensitive to the particular structure. This allows us to derive general conclusions on the effect of $z$, independently of the particular structure. For large $z$ quantum fluctuations are reduced and the specific heat approaches the mean-field value. Reducing $z$ the specific heat decreases below $T_{C}$ and increases above $T_{C}$ displaying larger tails.
For a quantitative comparison with experiments, it is necessary to consider also an axial anisotropy $H_{K}$ (see Section IV). The origin of $H_{K}$ is due to a small admixture of ${}^{6}P_{7/2}$ in the ground-state multiplet. The presence of a small component with $L=1$ renders the system sensitive to axial crystal fields.
Using $z$ as the only free parameter in addition to the critical temperature, we have fitted the specific heat of four Gd compounds (GdNi3Ga9, GdPdBi, GdCu2Ge2, and GdNiSi3). The specific heat of GdCoIn5 has been fitted previously using the same approach Facio et al. (2015). Curiously, we find that $H_{K}$ can be neglected for these Gd compounds. Including also $K$ as a parameter, we have done the same fitting for two Eu compounds (Eu2Pd3Sn3 and EuPdSn2). The agreement shown in Fig. 6 is noticeable. It is possible that the larger atomic volume of Eu+2 renders this ion more susceptible to crystal-field effects.
Note that while most of our calculations have been done on non-frustrated systems, some of the fitted compounds here (GdPdBi and GdNiSi3), and GdCoIn5 fitted previously Facio et al. (2015), are expected to have some degree of frustration.
Our work opens the possibility to classify and analyze the specific heat of non-frustrated Gd+3 and Eu+2, at least in a first semiquantitative manner using the above discussed two parameters.
@@ Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. AAA acknowledges financial support provided by PICT 2020A 03661 and PICT 2018-01546 of the Agencia I+D+i, Argentina. DJG is supported by PIP 11220200101796CO of CONICET, Argentina.

### v2

@@ IV Effect of anisotropy
In this Section, we address the role of anisotropy first in the mean-field approximation and then incorporate the effect of quantum fluctuations.
@@ IV.1 Mean-field Approximation
In this subsection, we discuss the effect of the term $H_{A}$ on the specific heat calculated in the mean-field approximation, in which it is assumed that each site is subjected to a constant effective magnetic field resulting from the interaction with the remaining sites. Specifically, the interactions appearing in $H_{0}$ [see Eq. (1)] are approximated as
$\mathbf{S}_{i}\cdot\mathbf{S}_{j}\approx\langle\mathbf{S}_{i}\rangle\cdot\mathbf{S}_{j}+\mathbf{S}_{i}\cdot\langle\mathbf{S}_{j}\rangle-\langle\mathbf{S}_{i}\rangle\cdot\langle\mathbf{S}_{j}\rangle,$ (4)
We assume that all sites are equivalent and there is no external magnetic field. For simplicity, we restrict the analysis to the easy axis case $K<0$ for which the spins point in the direction of the anisotropy axis $z$. In the next subsection, both signs of $K$ and the effects of quantum fluctuations will be considered. For an antiferromagnetic order, we rotate 180 degrees the spins pointing in the $-z$ direction around a perpendicular axis so that for all sites $\langle S_{z}\rangle\geq 0$. This expectation value serves as the order parameter for the magnetic phase.
@@
$\tilde{B}=g\mu_{B}B_{\text{eff}}=I\langle S_{z}\rangle\text{,~{} }\langle S_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{Me^{-\beta E_{M}}}{Z}\text{,~{}}Z=\sum\limits_{M=-J}^{J}e^{-\beta E_{M}},$ (5)
where $g$ is the gyromagnetic factor, $\mu_{B}$ is the Bohr magneton, $I=|\Sigma_{\delta}z_{\delta}J_{\delta}|$ is the sum of all interactions ($z_{\delta}$ is the number of neighbors at distance $\delta$ and $J_{\delta}$ its intensity), $\beta=1/(k_{B}T)$ where $T$ is the temperature and $k_{B}$ is the Boltzmann constant. The energies $E_{M}$ for each eigenvalue $M$ of the operator $S_{z}$ are given by:
$E_{M}=-\tilde{B}M+K[3M^{2}-S(S+1)].$ (6)
Eqs. (5) and (6) permit to determine $\tilde{B}$ self-consistently. In the limit $\tilde{B}\rightarrow 0$ one obtains an equation for the critical temperature
$k_{B}T_{C}^{\text{MF}}=I\langle S^{2}_{z}\rangle\text{, where~{}~{}}\langle S^{2}_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{M^{2}e^{-\beta E_{M}}}{Z}.$ (7)
The specific heat is obtained from the numerical derivative of the entropy $S$ with respect to temperature:
@@
In Fig. 3 we show the resulting effect of the axial anisotropy on the specific heat keeping the value of $k_{B}T_{C}=1$ constant, as the unit of energy. For $K=0$, the value of the intensity for a given $T_{C}$ is $I=3k_{B}T_{C}/[S(S+1)]$, leading to $I=4/21\sim 0.190$ for $k_{B}T_{C}=1$.
For $K=-0.02$, we obtain $I=0.1531$ from Eq. (7). This is because the expectation value $\langle S^{2}_{z}\rangle$ that enters Eq. (7) is larger for $K<0$ and should be compensated with a smaller $I$. In addition, since the values of high $S_{z}$ projection $|M|$ are favored, the effective magnetic field is more efficient than for $K=0$ in splitting the lowest energy levels and therefore the entropy and specific heat decrease for $T<0.4T_{C}$ with respect to the isotropic case. This is compensated by an increase in $C/T$ for $0.5T_{C}<T<T_{C}$.
@@
IV.2 Beyond Mean-field
In this subsection we discuss the effect of anisotropy in cases where the fluctuations of the effective field are also included.
Figure 4: (Color online) Specific heat over temperature as a function of temperature for several anisotropies for the simple cubic structure including 6 NN and 8 third NN with $J_{\delta}=-1$.
In Fig. 4 we show how the specific heat for a structure with $z=14$ is modified by the addition of $H_{A}$. The results are qualitatively the same as in the mean-field case for $T<T_{C}$. Addition of the anisotropy term shifts the entropy from the region $T<0.5T_{C}$ mainly to the region $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
However, in contrast to the mean-field case, addition of the anisotropy term decreases the specific heat for $T>T_{C}$. This reduction is expected because the specific heat in this region primarily results from quantum fluctuations, which are diminished by the energy level splitting induced by $H_{A}$.
@@ V Comparison of different structures with the same effective number of neighbors
In Section III, we discussed the effect of increasing the number of neighbors $z$ in the specific heat, while keeping $J_{\delta}=J_{1}$. Here we investigate the effect of changing $J_{\delta}$ for first, second, and possibly third nearest neighbors, keeping the same $r$ [see Eq. (3)], which corresponds to the inverse of the effective number of neighbors $z$. Some parameter sets may be highly unrealistic. However, the objective of this section is to demonstrate that entirely different parameter sets, which result in the same effective $z$, produce remarkably similar specific heat behaviors.
In Fig. 5 (b) we show the comparison for a moderately small $z$: a cubic BCC structure with 6 NN and 8 third NN. We compare the case of NN interaction $J_{1}=-1$ and NNN interaction $J_{2}=0$ with $J_{1}=-1/8$ and $J_{2}=-1$. In spite of the radical change in the magnitude of the interactions, the specific heats are similar.
In Fig. 5 (c), we show a similar comparison for an FCC structure with different combinations of $J_{1}$ and $J_{2}$. Again the results are similar, except for the case in which the dominant interaction is changed drastically from NN to NNN. When the dominant interaction is that of NNN, the specific heat is larger and the difference is of the order of that between different $z$ for $J_{1}=-1$ (Fig. 5 (a)). Finally, in Fig. 5 (d), we show a similar comparison for a larger $z$. In this case, the agreement between the two curves is excellent.
@@ VI Fits of experimental data
In this Section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. Previously, the specific heat of GdCoIn5 has been fitted [19].
The fitting parameters are the known $T_{C}$, which determines the magnitude of the interactions $J_{\delta}$ assumed equal, the number of neighbors $z$ and for the Eu compounds, the anisotropy $K$. The Gd compounds could be fitted with $K=0$. The assumption of $J_{\delta}$ independent of $\delta$ is justified from the results of the previous Section, which show that different $J_{\delta}$ with the same effective number of neighbors $1/r$ show very similar specific heats. The fact that $K$ is important only for the Eu compounds, might be expected for the $\sim 25\%$ larger ionic radius of Eu2+ compared to that of Gd3+ [28], rendering the former more sensitive to crystal-field effects.
The results are presented in Fig. 6. The Gd compounds are ordered with increasing effective number of neighbors $z$ in the fit. The parameters of the fits used for the different compounds are listed in Table 1. Also indicated is the number of NN in the reported structure, denoted as $z_{0}$. Notations like 1+1+2 in $z_{0}$ indicate that there are four atoms located at three nearly identical distances.
Compound $T_{N}$ (exp,[K]) $z_{0}$ $z$(fit) $J$[K] $K$[K]
GdNi3Ga9 20 3 4 1.52 0
GdCu2Ge2 12 4 8 0.38 0
GdPbBi 13 12 10 0.33 0
GdNiSi3 22 2+2+2 26 0.18 0
Eu2Pd3Sn 47 1+1+2 14 0.66 -0.26
EuPdSn2 12 2+2 14 0.55 -0.44
Table 1: Parameters used in the fits: $T_{N}$ is the Neel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure, $z$(fit) is the effective number of neighbors in the fit, $J$ is the intensity of the interactions, and $K$ is the anisotropy.
@@
GdNi3Ga9 [29, 30] clearly corresponds to a low coordination number ($z=4$). This result is expected, since the honeycomb lattice described in Section III with antiferromagnetic interactions (three intralayer and one interlayer for each site) corresponds to the structure of the material and provides a very good fit.
For the rest of the structures, ferromagnetic interactions are assumed. As stated in Section III, as long as there is no significant frustration, the difference between ferromagnetic and antiferromagnetic interactions for $z>6$ is only noticeable at very low temperatures, well below the temperature range of interest in this study. The key features of the specific heat curves we analyze, such as the shape of the peak and the high-temperature tail, are primarily determined by the effective coordination number and the single-ion anisotropy, and these features are largely insensitive to the sign of the exchange interaction in the absence of strong frustration.
@@
GdCu2Ge2 has a tetragonal structure near to a BCC one (Fig. 4.1 of Ref. [31]). Assuming superexchange interaction through Cu, the “BCC” like character is even more plausible. Using only 8 NN interactions of a BCC structure, the resulting specific heat is very near the experimental one, although slightly below (above) the experimental one for $T<T_{C}$ ($T>T_{C}$), suggesting that $z$ might be a little bit larger. If only the four NN were considered, the system would be two-dimensional (with a broadened transition near $T_{C}$) and this is incompatible with the observed specific heat.
GdPdBi [32] is a half-Heusler structure with antiferromagnetic coupling. We obtain a very good fit assuming a layered honeycomb lattice as for GdNi3Ga9, but with a larger number of effective neighbors (10). The fitting is better than using an FCC structure. This effect might be caused by a small distortion of the lattice. Note that in this case, the effective number of neighbors ($z=10$) is slightly lower than the number of nearest neighbors ($z_{0}=12$). This is likely due to a slight frustration, which enhances the relative impact of fluctuations.
@@
For GdNiSi3 [33], a simple cubic structure with $z=26$ including neighbors up to the third coordination shell provides a very good fit of the data. For a three-dimensional behavior, $z>16$ is necessary to connect different planes.
@@
This completes the Gd compounds that we have considered. In a previous work [19], the specific heat of GdCoIn5 has been fitted with essentially the same procedure. The compound has a tetragonal structure. The authors have extrapolated to the thermodynamic limit, the specific heat that corresponds to a simple cubic lattice $z=6$ NN with antiferromagnetic interactions, providing an excellent fit of the data.
@@
Now turning to Eu compounds, the specific heat of Eu2Pd3Sn [34] is very well fitted using a simple cubic structure with 6 NN and 8 third NN ($z=14$) with $J_{1}=J_{2}$ and $K=-0.4J_{1}$. Small deviations exist at very low temperatures, where our numerical method is not fully reliable. The point group at the Eu sites is $C_{1h}$, compatible with the presence of $H_{A}$ as discussed in the introduction.
@@
Finally, the same simple cubic structure, but with $K=-0.8J_{1}$ provides a very good fit of the specific heat of EuPdSn2 [35]. For temperatures very near $T_{C}$, the theoretical data are a little above (below) the experimental ones for $T<T_{C}$ ($T>T_{C}$). The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. The discrepancies near $T_{C}$ ($T_{N}$) likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 15.
Since all the compounds mentioned above contain transition-metal ions, which in some cases are magnetic, one might question whether they influence the reported specific heat. However, our analysis shows that the entropy obtained by integrating $C/T$ is very near ln(8) per rare-earth ion (consistent with $S=7/2$). This indicates that any effect from the transition metals is negligible.
@@
Figure 6: (Color online) Specific heat over temperature as a function of temperature for several compounds and our fit. $K=0$ for the Gd compounds. Experimental data were taken from Refs. 29, 30, 32, 31, 33, 35, 34.
@@ VII Summary
Using quantum Monte Carlo, we have calculated the specific heat of a large number of non-frustrated magnetic structures. As expected from mean-field theory, the critical temperature $T_{C}$ increases approximately proportionally to $I=\Sigma_{\delta}z_{\delta}|J_{\delta}|$, where $J_{\delta}$ is the exchange interaction at distance $\delta$ and $z_{\delta}$ is the number of neighbors at this distance. However, $T_{C}$ lies below the mean-field value.
Rather surprisingly, we find that the specific heat depends mainly on the effective number of neighbors $z$ [see Eq. (3)] and is not quite sensitive to the particular structure. This allows us to derive general conclusions on the effect of $z$, independently of the particular structure. For large $z$, quantum fluctuations are reduced, and the specific heat approaches the mean-field value. Reducing $z$, the specific heat decreases below $T_{C}$ and increases above $T_{C}$ displaying larger tails.
For a quantitative comparison with experimental results, it is essential to incorporate axial anisotropy $H_{A}$ (see Section IV). The origin of $H_{A}$ lies in a small admixture of ${}^{6}P_{7/2}$ in the ground-state multiplet. This admixture introduces a component with $L=1$, rendering the system sensitive to axial crystal fields.
Using $z$ as the only free parameter in addition to the critical temperature, we have fitted the specific heat of four Gd compounds (GdNi3Ga9, GdPdBi, GdCu2Ge2, and GdNiSi3). The specific heat of GdCoIn5 has been fitted previously using the same approach [19]. Curiously, we find that $H_{A}$ can be neglected for these Gd compounds. Including also $K$ as a parameter, we have done the same fitting for two Eu compounds (Eu2Pd3Sn3 and EuPdSn2). The agreement shown in Fig. 6 is noticeable. It is possible that the larger atomic volume of Eu+2 renders this ion more susceptible to crystal-field effects.
Note that while most of our calculations have been done on non-frustrated systems, some of the fitted compounds here (GdPdBi and GdNiSi3), and GdCoIn5 fitted previously [19], are expected to have some degree of frustration.
Our work opens the possibility to classify and analyze the specific heat of non-frustrated Gd+3 and Eu+2, at least in a first semiquantitative manner using the above-discussed two parameters.
@@ Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. A.A.A. acknowledges financial support provided by PICT 2020A 03661 of the Agencia I+D+i, Argentina. D.J.G. is supported by PIP 11220200101796CO of CONICET, Argentina.